OpenStudy (anonymous):

Write an equation of the line perpendicular to the given line that contains P. y space equals space 1 fifth x space plus space 1; P(2, -7)

OpenStudy (anonymous):

help me pleaase!!!

OpenStudy (irishboy123):

lmLtfy! "y space equals space 1 fifth x space plus space 1" is: \(y = \frac{1}{5}x+ 1\) and hint: if the lines \(y = mx + c\) and \(y = m' x + c'\) are perp, then \(m' = -\frac{1}{m}\) and \(m = -\frac{1}{m'}\)

OpenStudy (anonymous):

im osrry i still dont understand

OpenStudy (irishboy123):

what is the gradient of your line?

OpenStudy (anonymous):

there is no line

OpenStudy (anonymous):

is the answer y=-1/5x-33/5??

OpenStudy (irishboy123):

go back to your question you have a line

OpenStudy (irishboy123):

i did the Latex for you

OpenStudy (irishboy123):

no m = -1 / m' m is slope of original line m' is slope of line that is perp to original line

OpenStudy (anonymous):

oh right so then would it be y=1/5x+3

OpenStudy (irishboy123):

i will keep asking this :p what is the slope of your original line, the one that you posted??

OpenStudy (irishboy123):

do you know what a slope is? \(\frac{rise}{run}\) ??

OpenStudy (anonymous):

1/5?

OpenStudy (irishboy123):

great so m' = -1 / m m is slope of original line m' is slope of line that is perp to original line

OpenStudy (anonymous):

so then it would be -1/5?

OpenStudy (irishboy123):



OpenStudy (irishboy123):

you're just "minusing" it. you need to invert and minus it